OpenStudy (anonymous):

Find the radius of a circle whose area equals the area of a rectangle that measures 2 ft. by 11 ft. (Use = 22/7)

OpenStudy (anonymous):

\[ 2\times 11 = \frac {22} 7 \times r^2 \]

jhonyy9 (jhonyy9):

so hope you know formula for area of a circle and formula for area of a rectangle than write it and make is equale ...and from these you will can calculi the radius

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."